# Camera-based 3D Semantic Scene Completion with Sparse Guidance Network

### v1

@@ Introduction
Figure 3: Detailed architecture of the proposed semantic guidance module. Sparse encoder block (SEB) consists of a sparse feature encoder and a sparse geometry feature encoder adopted from Ye et al. (2022).
Sparse Semantic Guidance.
Since directly learning the semantics of all the voxels from the 3D features with coarse volumetric information is less effective and efficient, we propose propagating semantics from seed voxel to the whole scene. Specifically, we generate sparse voxel proposals to choose seed voxels and encourage inter-category separability of seed features with semantic guidance, expediting the latter semantic diffusion.
Sparse Voxel Proposal.
We devise the sparse voxel proposal network (SVPN) to generate sparse proposals for indexing seed voxels. Unlike Voxformer Li et al. (2023d), which learns class-agnostic proposals on temporal data offline for voxel queries, our SVPN aims to dynamically select seed voxels online by occupancy probability for subsequent semantic context learning. Specifically, SVPN consists of depth estimation and coarse-to-fine occupancy prediction. Following Li et al. (2023d); Jiang et al. (2023), we utilize the pre-trained Mobilestereonet Shamsafar et al. (2022) to infer the depth prediction and calculate the scene points P by back-projecting the depth map into the 3D point cloud using the camera parameters $(K,T)$. The scene points P imply the volumetric surface and are used for occupancy prediction.
Next, we generate the occupancy prediction $\textbf{O}\in\mathbb{R}^{X\times Y\times Z}$ in a coarse-to-fine manner. Firstly, the points P are fed into a voxelization layer adopted from DRNet Ye et al. (2021) for voxel-wise features. Then we apply the tiny sparse convolution network consisting of a sparse feature encoder and a sparse geometry feature encoder adopted from GASN Ye et al. (2022) to predict the coarse occupancy probability from the voxel-wise features. Finally, the occupancy probability is further fed into a lightweight Unet-like network Roldão, de Charette, and Verroust-Blondet (2020) for the final occupancy prediction O, which is used to select the sparse voxels as explained in semantic guidance. Similar to the geometry guidance, we use the binary cross entropy loss to calculate the loss $\mathcal{L}_{occ}$ for occupancy prediction.
To utilize the occupancy-aware cues, we also take the 3D features $\textbf{F}_{o}^{3D}\in\mathbb{R}^{C_{o}\times X\times Y\times Z}$ from the last layer of the Unet-like network for the latter voxel feature aggregation.
Semantic Guidance.
After obtaining the occupancy prediction O and voxel coordinates $V^{3D}\in\mathbb{Z}^{3\times X\times Y\times Z}$ of the scene, we first choose the initial seed voxel features $\textbf{F}_{s,0}^{3D}\in\mathbb{R}^{C\times N_{s}}$ and seed coordinates $V_{s}^{3D}\in\mathbb{Z}^{3\times N_{s}}$ by:
$V_{s}^{3D}=V^{3D}[:,\textbf{O}>\theta]$ (4)
@@
where $\theta$ is the threshold to determine if the voxel is occupied and $N_{s}$ is the number of non-empty voxel. Then these seed voxel features $\textbf{F}_{s,0}^{3D}$ and corresponding voxel indices $V_{s}^{3D}$ are fed into the semantic guidance module illustrated in Fig.3 for mutual interactions. The semantic guidance module has two sparse encoder blocks (SEB), a fusion layer, and an auxiliary semantic head. Each encoder block consists of a sparse feature encoder and a sparse geometry feature encoder adopted from Ye et al. (2022) and outputs features with multi-scale contextual information. Let $\textbf{F}_{s,1}^{3D},\textbf{F}_{s,2}^{3D}$ are the outputs of the two sparse encoder blocks, the fusion feature $\textbf{F}_{s}^{3D}\in\mathbb{R}^{C\times N_{s}}$ are obtained by:
@@
where [.][.][ . ] denotes concatenate operation along feature dimension. After that, the fused features $\textbf{F}_{s}^{3D}$ are fed into the auxiliary semantic head consisting of a two-layer MLP to predict the corresponding semantics $\hat{\textbf{Y}}_{s}\in\mathbb{R}^{C_{class}\times N_{s}}$, where $C_{class}$ is the number of classes. We calculate the cross entropy loss and lovasz loss Berman, Rannen Triki, and Blaschko (2018) for the semantic guidance:
@@
By this means, we inject semantic cues into the fused seed features $\textbf{F}_{s}^{3D}$ and enhance the feature separation between categories, which is the key to semantic diffusion.
@@
${}^{\dagger}$ Roldão, de Charette, and Verroust-Blondet (2020) 31.38 46.70 19.50 13.50 3.10 10.30 14.30 0.30 0.00 0.00 0.00 10.80 0.00 10.40 0.00 0.00 0.00 5.40 0.00 0.00 7.07
${}^{\dagger}$ Li et al. (2020) 23.93 39.30 18.30 19.80 1.60 9.60 15.30 0.70 0.00 0.00 0.00 9.60 1.90 13.50 0.00 0.00 0.00 5.00 0.10 0.00 7.09
${}^{\dagger}$ Yan et al. (2021) 34.00 47.30 21.70 19.90 2.80 12.70 20.10 0.80 0.00 0.00 4.10 14.20 3.10 12.40 0.00 0.20 0.20 8.70 1.90 0.30 8.97
MonoScene Cao and de Charette (2022) 34.16 54.70 27.10 24.80 5.70 14.40 18.80 3.30 0.50 0.70 4.40 14.90 2.40 19.50 1.00 1.40 0.40 11.10 3.30 2.10 11.08
TPVFormer Huang et al. (2023) 34.25 55.10 27.20 27.40 6.50 14.80 19.20 3.70 1.00 0.50 2.30 13.90 2.60 20.40 1.10 2.40 0.30 11.00 2.90 1.50 11.26
VoxFormer Li et al. (2023d) 42.95 53.90 25.30 21.10 5.60 19.80 20.80 3.50 1.00 0.70 3.70 22.40 7.50 21.30 1.40 2.60 0.20 11.10 5.10 4.90 12.20
OccFormer Zhang, Zhu, and Du (2023) 34.53 55.90 30.30 31.50 6.50 15.70 21.60 1.20 1.50 1.70 3.20 16.80 3.90 21.30 2.20 1.10 0.20 11.90 3.80 3.70 12.32
SurroundOcc Wei et al. (2023) 34.72 56.90 28.30 30.20 6.80 15.20 20.60 1.40 1.60 1.20 4.40 14.90 3.40 19.30 1.40 2.00 0.10 11.30 3.90 2.40 11.86
Symphonize Jiang et al. (2023) 41.07 55.70 26.80 25.30 4.90 21.30 22.10 1.90 1.70 1.30 5.80 22.90 8.20 19.50 2.20 1.30 0.50 13.10 6.80 5.80 13.02
@@
Table 1: Semantic Scene Completion on SemanticKITTI hidden test set. ${}^{\dagger}$ denotes the results provided by MonoScene Cao and de Charette (2022). Bold and underline denote the best performance and the second-best performance, respectively.
@@
${}^{\dagger}$ Roldão, de Charette, and Verroust-Blondet (2020) 28.61 40.68 18.22 4.38 0.00 10.31 18.33 0.00 0.00 0.00 0.00 13.66 0.02 20.54 0.00 0.00 0.00 1.21 0.00 0.00 6.70
${}^{\dagger}$ Li et al. (2020) 29.59 43.55 20.55 11.97 0.07 12.94 14.71 4.53 0.00 0.00 0.00 15.37 2.90 28.71 0.00 0.00 0.00 2.52 0.06 0.00 8.31
${}^{\dagger}$ Yan et al. (2021) 38.98 50.49 23.74 11.94 0.07 15.03 24.65 4.41 0.00 0.00 6.15 18.11 4.33 26.86 0.67 0.27 0.00 3.94 3.77 1.45 10.31
MonoScene Cao and de Charette (2022) 37.12 57.47 27.05 15.72 0.87 14.24 23.55 7.83 0.20 0.77 3.59 18.12 2.57 30.76 1.79 1.03 0.00 6.39 4.11 2.48 11.50
TPVFormer Huang et al. (2023) 35.61 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52 11.36
VoxFormer Li et al. (2023d) 44.02 54.76 26.35 15.50 0.70 17.65 25.79 5.63 0.59 0.51 3.77 24.39 5.08 29.96 1.78 3.32 0.00 7.64 7.11 4.18 12.35
OccFormer Zhang, Zhu, and Du (2023) 36.50 58.85 26.88 19.61 0.31 14.40 25.09 25.53 0.81 1.19 8.52 19.63 3.93 32.62 2.78 2.82 0.00 5.61 4.26 2.86 13.46
Symphonize Jiang et al. (2023) 41.44 55.78 26.77 14.57 0.19 18.76 27.23 15.99 1.44 2.28 9.52 24.50 4.32 28.49 3.19 8.09 0.00 6.18 8.99 5.39 13.44
@@
Table 2: Semantic Scene Completion on SemanticKITTI val set. ${}^{\dagger}$ denotes the results provided by MonoScene.
Voxel Aggregation.
@@
As shown in Fig.2, we further aggregate the semantic-aware seed features $\textbf{F}_{s}^{3D}$ and occupancy-aware features $\textbf{F}_{o}^{3D}$ with 3D features $\textbf{F}^{3D}$ implying coarse geometry representation to construct the final discriminative voxel features $\textbf{F}_{f}^{3D}\in\mathbb{R}^{(C+C_{o})\times X\times Y\times Z}$. Specifically, we leverage the coordinates of non-seed voxels to index features $\textbf{F}_{n}^{3D}$ from $\textbf{F}^{3D}$. Then the non-seed voxel features $\textbf{F}_{n}^{3D}$ are fed into a knowledge transfer layer and combined with the semantic-aware features $\textbf{F}_{s}^{3D}$ to form the new scene representation, which contains the semantic context and geometry prior. We argue that non-seed voxel features $\textbf{F}_{n}^{3D}$ are vital and can well complement the seed features since the sparse voxel proposals are not always accurate. To exploit spatial occupancy information from the SVPN, we also concatenate the occupancy-aware features $\textbf{F}_{o}^{3D}$ with the new scene representation to obtain the final voxel features. The detailed procedure can be formulated as follows:
$\textbf{F}_{f}^{3D}=\mathrm{MLP}([\mathrm{CN}(\textbf{F}_{s}^{3D},\mathrm{KT}(\textbf{F}_{n}^{3D})),\textbf{F}_{o}^{3D}])$ (8)
where $\mathrm{CN}$ is the feature combination of seed and non-seed voxels, and $\mathrm{KT}$ is the linear layer for knowledge transfer.
Multi-Scale Semantic Diffusion

### v2

@@ III-B Feature Learning with Hybrid Guidance
Figure 3: Detailed architecture of the proposed semantic guidance module (SGM). The sparse encoder block (SEB) consists of a sparse feature encoder and a sparse geometry feature encoder adopted from [52].
III-B2 Sparse Semantic Guidance.
Since directly learning the semantics of all the voxels from the 3D features with coarse volumetric information is less effective and efficient, we propose propagating semantics from seed voxel to the whole scene. Specifically, we generate the sparse voxel proposal to choose seed voxels and encourage inter-category separability of seed features with semantic guidance, expediting the latter semantic propagation.
Sparse Voxel Proposal
[…]
$p=\textbf{R}^{-1}\cdot[\textbf{K}^{-1}\cdot(z_{c}\circ[u,v,1]^{T})-\textbf{t}]$ (6)
where $\circ$ denote element-wise product, $z_{c}$ is the depth value of the pixel. K and $\textbf{T}=[\textbf{R},\textbf{t}]$ are the intrinsic and extrinsic parameters of the camera.
Next, we generate the occupancy prediction $\textbf{O}\in\mathbb{R}^{X\times Y\times Z}$ in a coarse-to-fine manner. Firstly, the points P are fed into a voxelization layer adopted from DRNet [54] for voxel-wise features. Then we apply the tiny sparse convolution network consisting of a sparse feature encoder and a sparse geometry feature encoder adopted from GASN [52] to predict the coarse occupancy probability from the voxel-wise features. Finally, the occupancy probability is further fed into a lightweight Unet-like network [1] for the final occupancy prediction O, which is used to select the sparse voxels as explained in semantic guidance. Similar to the geometry guidance, we use the binary cross entropy loss to calculate the loss $\mathcal{L}_{occ}$ for occupancy prediction.
To further utilize the geometry information from the depth-based SVPN, we also take the occupancy-aware 3D features $\textbf{F}_{o}^{3D}\in\mathbb{R}^{C_{o}\times X\times Y\times Z}$ from the last layer of the Unet-like network for the latter voxel feature aggregation.
Semantic Guidance
After obtaining the occupancy prediction O and voxel coordinates $\textbf{V}^{3D}\in\mathbb{Z}^{3\times X\times Y\times Z}$ of the scene, we first choose the initial seed voxel features $\textbf{F}_{s,0}^{3D}\in\mathbb{R}^{C\times N_{s}}$ and seed coordinates $\textbf{V}_{s}^{3D}\in\mathbb{Z}^{3\times N_{s}}$ by:
$\textbf{V}_{s}^{3D}=\textbf{V}^{3D}[:,\textbf{O}>\theta]$ (7)
@@
where $\theta$ is the threshold to determine if the voxel is occupied and $N_{s}$ is the number of non-empty voxel. Then these seed voxel features $\textbf{F}_{s,0}^{3D}$ and corresponding voxel indices $V_{s}^{3D}$ are fed into the semantic guidance module (SGM) illustrated in Figure 3 for mutual interactions. The semantic guidance module has two sparse encoder blocks (SEB), a fusion layer, and an auxiliary semantic head. Each sparse encoder block (SEB) consists of a sparse feature encoder and a sparse geometry feature encoder adopted from [52] and outputs features with multi-scale contextual information. Let $\textbf{F}_{s,1}^{3D},\textbf{F}_{s,2}^{3D}$ are the outputs of the two sparse encoder blocks, the fusion feature $\textbf{F}_{s}^{3D}\in\mathbb{R}^{C\times N_{s}}$ are obtained by:
@@
where [.][.][ . ] denotes concatenate operation along feature dimension. After that, the fused features $\textbf{F}_{s}^{3D}$ are fed into the auxiliary semantic head consisting of a two-layer MLP to predict the corresponding semantics $\hat{\textbf{Y}}_{s}\in\mathbb{R}^{C_{class}\times N_{s}}$, where $C_{class}$ is the number of classes. We calculate the cross entropy loss and lovasz loss [55] for the semantic guidance:
@@
By this means, we inject semantic cues into the fused seed features $\textbf{F}_{s}^{3D}$ and enhance the feature separation between categories, which is the key to semantic propagation.
TABLE I: Semantic Scene Completion on SemanticKITTI hidden test set. denotes the results provided by MonoScene [7]. Bold and underline denote the best performance and the second-best performance, respectively.
@@
LMSCNet [1] 31.38 46.70 19.50 13.50 3.10 10.30 14.30 0.30 0.00 0.00 0.00 10.80 0.00 10.40 0.00 0.00 0.00 5.40 0.00 0.00 7.07
AICNet [13] 23.93 39.30 18.30 19.80 1.60 9.60 15.30 0.70 0.00 0.00 0.00 9.60 1.90 13.50 0.00 0.00 0.00 5.00 0.10 0.00 7.09
JS3C-Net [4] 34.00 47.30 21.70 19.90 2.80 12.70 20.10 0.80 0.00 0.00 4.10 14.20 3.10 12.40 0.00 0.20 0.20 8.70 1.90 0.30 8.97
MonoScene [7] 34.16 54.70 27.10 24.80 5.70 14.40 18.80 3.30 0.50 0.70 4.40 14.90 2.40 19.50 1.00 1.40 0.40 11.10 3.30 2.10 11.08
TPVFormer [29] 34.25 55.10 27.20 27.40 6.50 14.80 19.20 3.70 1.00 0.50 2.30 13.90 2.60 20.40 1.10 2.40 0.30 11.00 2.90 1.50 11.26
VoxFormer [12] 42.95 53.90 25.30 21.10 5.60 19.80 20.80 3.50 1.00 0.70 3.70 22.40 7.50 21.30 1.40 2.60 0.20 11.10 5.10 4.90 12.20
OccFormer [10] 34.53 55.90 30.30 31.50 6.50 15.70 21.60 1.20 1.50 1.70 3.20 16.80 3.90 21.30 2.20 1.10 0.20 11.90 3.80 3.70 12.32
SurroundOcc [9] 34.72 56.90 28.30 30.20 6.80 15.20 20.60 1.40 1.60 1.20 4.40 14.90 3.40 19.30 1.40 2.00 0.10 11.30 3.90 2.40 11.86
NDC-scene [32] 36.19 58.12 28.05 25.31 6.53 14.90 19.13 4.77 1.93 2.07 6.69 17.94 3.49 25.01 3.44 2.77 1.64 12.85 4.43 2.96 12.58
@@
TABLE II: Semantic Scene Completion on SemanticKITTI val set. denotes the results provided by MonoScene. Bold and underline denote the best performance and the second-best performance, respectively.
@@
LMSCNet [1] 28.61 40.68 18.22 4.38 0.00 10.31 18.33 0.00 0.00 0.00 0.00 13.66 0.02 20.54 0.00 0.00 0.00 1.21 0.00 0.00 6.70
AICNet [13] 29.59 43.55 20.55 11.97 0.07 12.94 14.71 4.53 0.00 0.00 0.00 15.37 2.90 28.71 0.00 0.00 0.00 2.52 0.06 0.00 8.31
JS3C-Net [4] 38.98 50.49 23.74 11.94 0.07 15.03 24.65 4.41 0.00 0.00 6.15 18.11 4.33 26.86 0.67 0.27 0.00 3.94 3.77 1.45 10.31
MonoScene [7] 37.12 57.47 27.05 15.72 0.87 14.24 23.55 7.83 0.20 0.77 3.59 18.12 2.57 30.76 1.79 1.03 0.00 6.39 4.11 2.48 11.50
TPVFormer [29] 35.61 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52 11.36
VoxFormer [12] 44.02 54.76 26.35 15.50 0.70 17.65 25.79 5.63 0.59 0.51 3.77 24.39 5.08 29.96 1.78 3.32 0.00 7.64 7.11 4.18 12.35
OccFormer [10] 36.50 58.85 26.88 19.61 0.31 14.40 25.09 25.53 0.81 1.19 8.52 19.63 3.93 32.62 2.78 2.82 0.00 5.61 4.26 2.86 13.46
NDC-scene [32] 37.24 59.20 28.24 21.42 1.67 14.94 26.26 14.75 1.67 2.37 7.73 19.09 3.51 31.04 3.60 2.74 0.00 6.65 4.53 2.73 12.70
@@
TABLE III: Quantitative comparison in different ranges on SemanticKITTI validation. “*” denotes the results provided by VoxFormer.
Methods Modality IoU (%) mIoU (%)
12.8m 25.6m 51.2m 12.8m 25.6m 51.2m
SSCNet [16] LiDAR 64.37 61.02 50.22 20.02 19.68 16.35
JS3CNet [4] LiDAR 63.47 63.40 53.09 30.55 28.12 22.67
MonoScene* [7] Camera 38.42 38.55 36.80 12.25 12.22 11.30
OccFormer [10] Camera 56.38 47.28 36.50 20.91 17.90 13.46
VoxFormer-S [12] Camera 65.35 57.54 44.02 17.66 16.48 12.35
VoxFormer-T [12] Camera 65.38 57.69 44.15 21.55 18.42 13.35
SGN-S (ours) Camera 64.21 56.20 43.60 21.53 19.60 14.55
SGN-L (ours) Camera 70.08 61.17 45.45 24.76 21.17 14.80
SGN-T (ours) Camera 70.61 61.90 46.21 25.70 22.02 15.32
TABLE IV: Quantitative results on SSCBench-KITTI360 test set. The results for counterparts are provided in [15]. Bold and underline denote the best performance and the second-best performance, respectively.
Method IoU Precision Recall mIoU
$\blacksquare$ car (2.85%)
$\blacksquare$ bicycle (0.01%)
$\blacksquare$ motorcycle (0.01%)
$\blacksquare$ truck (0.16%)
$\blacksquare$ other-veh. (5.75%)
$\blacksquare$ person (0.02%)
$\blacksquare$ road (14.98%)
$\blacksquare$ parking (2.31%)
$\blacksquare$ sidewalk (6.43%)
$\blacksquare$ other-grnd. (2.05%)
$\blacksquare$ building (15.67%)
$\blacksquare$ fence (0.96%)
$\blacksquare$ vegetation (41.99%)
$\blacksquare$ terrain (7.10%)
$\blacksquare$ pole (0.22%)
$\blacksquare$ traf.-sign (0.06%)
$\blacksquare$ other-struct. (4.33%)
$\blacksquare$ other-obj. (0.28%)
LiDAR-based
SSCNet [16] 53.58 69.63 69.92 16.95 31.95 0.00 0.17 10.29 0.00 0.07 65.70 17.33 41.24 3.22 44.41 6.77 43.72 28.87 0.78 0.75 8.69 0.67
LMSCNet [1] 47.35 72.77 57.55 13.65 20.91 0.00 0.00 0.26 0.58 0.00 62.95 13.51 33.51 0.20 43.67 0.33 40.01 26.80 0.00 0.00 3.63 0.00
Camera-based
MonoScene [7] 37.87 56.73 53.26 12.31 19.34 0.43 0.58 8.02 2.03 0.86 48.35 11.38 28.13 3.32 32.89 3.53 26.15 16.75 6.92 5.67 4.20 3.09
TPVFormer [29] 40.22 59.32 55.54 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
VoxFormer [12] 38.76 58.52 53.44 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
OccFormer [10] 40.27 59.70 55.31 13.81 22.58 0.66 0.26 9.89 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
DepthSSC [56] 40.85 60.69 55.86 14.28 21.90 2.36 4.30 11.51 4.56 2.92 50.88 12.89 30.27 2.49 37.33 5.22 29.61 21.59 5.97 7.71 5.24 3.51
Symphonize [30] 44.12 69.24 54.88 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
SGN-S (ours) 46.22 68.17 58.94 17.71 28.20 2.09 3.02 11.95 3.68 4.20 59.49 14.50 36.53 4.24 39.79 7.14 36.61 23.10 14.86 16.14 8.24 4.95
SGN-L (ours) 46.64 68.26 59.55 16.95 26.91 1.72 0.85 8.60 3.80 1.93 56.52 13.83 35.40 3.42 40.62 6.65 36.68 22.00 15.84 16.49 8.06 5.76
SGN-T (ours) 47.06 68.83 59.81 18.25 29.03 3.43 2.90 10.89 5.20 2.99 58.14 15.04 36.40 4.43 42.02 7.72 38.17 23.22 16.73 16.38 9.93 5.86
III-C Voxel Aggregation
@@
As shown in Figure 2, to fully exploit the geometry information in 3D features $\textbf{F}^{3D}$ and $\textbf{F}_{o}^{3D}$, we further aggregate them with the semantic-aware seed features $\textbf{F}_{s}^{3D}$ to construct the final discriminative voxel features $\textbf{F}_{f}^{3D}\in\mathbb{R}^{(C+C_{o})\times X\times Y\times Z}$ for subsequent semantic propagation. Specifically, we leverage the coordinates of non-seed voxels to index features $\textbf{F}_{n}^{3D}$ from $\textbf{F}^{3D}$. Then the non-seed voxel features $\textbf{F}_{n}^{3D}$ are fed into a linear layer and combined with the semantic-aware features $\textbf{F}_{s}^{3D}$ to form the new scene representation, which contains the semantic context and geometry cues. We argue that non-seed voxel features $\textbf{F}_{n}^{3D}$ are vital and can well complement the seed features since the sparse voxel proposal is not always accurate. To further utilize the geometry information from the SVPN, we also concatenate the features $\textbf{F}_{o}^{3D}$ from SVPN with the new scene representation to obtain the final voxel features. The detailed procedure can be formulated as follows:
$\textbf{F}_{f}^{3D}=\mathrm{MLP}([\mathrm{CN}(\textbf{F}_{s}^{3D},\mathrm{Conv1d}(\textbf{F}_{n}^{3D})),\textbf{F}_{o}^{3D}])$ (11)
where $\mathrm{CN}$ is the feature combination of seed and non-seed voxels.
III-D Multi-Scale Semantic Propagation